### v2

@@ Abstract
The limited availability of labeled data is a major challenge in audiovisual speech emotion recognition (SER). Self-supervised learning approaches have recently been proposed to mitigate the need for labeled data in various applications. This paper proposes the VQ-MAE-AV model, a vector quantized masked autoencoder (MAE) designed for audiovisual speech self-supervised representation learning and applied to SER. Unlike previous approaches, the proposed method employs a self-supervised paradigm based on discrete audio and visual speech representations learned by vector quantized variational autoencoders. A multimodal MAE with self- or cross-attention mechanisms is proposed to fuse the audio and visual speech modalities and to learn local and global representations of the audiovisual speech sequence, which are then used for an SER downstream task. Experimental results show that the proposed approach, which is pre-trained on the VoxCeleb2 database and fine-tuned on standard emotional audiovisual speech datasets, outperforms the state-of-the-art audiovisual SER methods. Extensive ablation experiments are also provided to assess the contribution of the different model components.
@@
Self-supervised learning , Masked autoencoder , Audiovisual speech processing, Emotion recognition
@@ 1 Introduction
Emotion recognition [El Ayadi et al., 2011] aims to automatically identify emotions from unimodal or multimodal data. With the increasing prevalence of audiovisual interfaces, there is a growing need for systems that can accurately recognize emotions from audiovisual speech data. The complementarity of the audio and visual modalities has been demonstrated in various tasks [Gan et al., 2020, Gao & Grauman, 2019, Zhao et al., 2019, Sadok et al., 2024], and using them together can lead to more effective emotion recognition systems. According to Mehrabian [2017], only a small proportion of feelings and emotions is conveyed through the words we use in oral communication. The majority of our feelings and emotions is conveyed through tone, voice, facial expressions, and body language. In recent years, advances in artificial intelligence technology and hardware acceleration have enabled a shift towards multimodal processing in the field of emotion recognition [Ramachandram & Taylor, 2017, Tsai et al., 2019, Schoneveld et al., 2021]. Supervised learning using large annotated datasets remains the dominating paradigm. However, collecting a substantial amount of emotionally labeled data can be expensive, time-consuming, and sometimes impractical. Most datasets for emotion recognition rely on actors to perform various emotions at specific intensities, which requires a considerable amount of acquisition time and resources. In addition, in-the-wild datasets suffer from annotation issues, where annotators fail to reach a consensus, resulting in a significant amount of unusable data [Busso et al., 2008]. Unsupervised and self-supervised learning (SSL) approaches are natural ways to address these issues and have been explored in recent studies [Bengio et al., 2006, Wang et al., 2021, Dib et al., 2023],[Chen & Rudnicky, 2023]. SSL offers the advantage of high scalability, as the SSL task can be carried out on large amounts of unlabeled speech data [Liu et al., 2022, Zhang et al., 2022]. These methods involve pre-training models with pretext tasks and then fine-tuning them on a smaller set of labeled data for the emotion recognition task [Gong et al., 2022a, Pepino et al., 2021, Jegorova et al., 2023].
@@
Figure 1: Discrete audio and visual tokens creation: (i) fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1); (ii) discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2).
In this paper, we introduce the VQ-MAE-AV model, a vector quantized (VQ) masked autoencoder (MAE) for audiovisual (AV) speech representation learning applied to audiovisual SER. The overall architecture of the model is illustrated in Fig. 1 and 2. To the best of our knowledge, this is the first multimodal SSL approach based on MAEs [He et al., 2022] that is proposed for audiovisual SER. Moreover, unlike existing multimodal MAEs proposed in other application contexts [Bachmann et al., 2022, Geng et al., 2022], the VQ-MAE-AV model operates on discrete latent representations of the modalities and introduces global tokens that are learned using contrastive learning to capture the overall sequence-wise information from each modality. The VQ-MAE-AV model is pre-trained on the VoxCeleb2 dataset [Chung et al., 2018] and fine-tuned on two standard emotional audiovisual speech datasets. The experimental results demonstrate that, across the two test datasets, the proposed VQ-MAE-AV model outperforms the state-of-the-art audiovisual SER methods. Extensive ablation experiments are also presented to study the impact of different model designs. The code and qualitative results are available at https://samsad35.github.io/VQ-MAE-AV.
Figure 2: VQ-MAE-AV model structure. See the first paragraph of Section 3 for a complete description of the pipeline.
@@ 2 Related work
SSL models can be broadly categorized into discriminative and generative approaches [Zhang et al., 2022]. Discriminative SSL focuses on creating pairs or groups of data samples and formulating loss functions that enable the model to distinguish or group these samples, which can later benefit downstream tasks [Chen et al., 2020b, a]. Pretext tasks can consist of solving jigsaw puzzles [Noroozi & Favaro, 2016] or predicting image rotations Gidaris et al. [2018] for example. Contrastive learning has emerged as the predominant paradigm in discriminative SSL [Alayrac et al., 2020, Akbari et al., 2021]. In contrast, generative SSL involves generating or reconstructing segments of unlabeled and potentially corrupted data using an autoencoder model [He et al., 2022, Bao et al., 2021, Xie et al., 2022]. The latent representation produced by the encoder can subsequently be used for downstream tasks.
@@
The present paper focuses on the MAE, an SSL generative model that uses an asymmetric encoder-decoder architecture with input masking [He et al., 2022]. The MAE approach is inspired by masked language modeling [Devlin et al., 2019] and has been successfully applied to image modeling thanks to the development of Vision Transformers (ViT) [Dosovitskiy et al., 2020]. The MAE has recently been adapted for audio using a 2D time-frequency representation [Gong et al., 2022a, Baade et al., 2022, Huang et al., 2022]. In the MAE paradigm, the input is divided into non-overlapping patches, each represented by a token embedding. A proportion of tokens are masked, typically 75% for image/audio modeling and 15% for text modeling, and only the visible tokens are fed to the encoder. The encoder outputs are used to reconstruct the masked tokens through a lightweight decoder. To successfully reconstruct the masked tokens, the encoder must generate a semantic representation, which can then be effectively transferred to downstream tasks [He et al., 2022]. It was recently shown that combining the task of reconstructing masked tokens with contrastive learning can improve the representation learned by an MAE [Huang et al., 2023, Gong et al., 2022a].
Extensions of MAE for sequential and multimodal representation learning
A recent extension of the MAE was presented in [Feichtenhofer et al., 2022, Tong et al., 2022] for modeling image sequences, called Video-MAE. It uses the same architecture as the vanilla MAE [He et al., 2022] but incorporates a masking process from video ViT (ViViT) [Arnab et al., 2021]. Since videos often contain redundant information, particularly in scenes with no motion, the authors propose a cubic masking approach along the temporal axis, combined with a high masking ratio of 90%. Other works have extended the MAE to handle multimodal data [Bachmann et al., 2022, Geng et al., 2022, Gong et al., 2022b]. MultiMAE [Bachmann et al., 2022] encodes a small random subset of visible tokens from multiple modalities (RGB, depth, and semantic images) and is trained to reconstruct the missing ones. M3AE [Geng et al., 2022] is a unified MAE architecture for two input modalities (image and text). The main difference in architecture between M3AE and MultiMAE lies in the architecture of the decoder. The MulitMAE approach introduces individual decoders for each modality, enhancing their fusion by incorporating a cross-attention layer at the outset of each decoder. On the other hand, the M3AE approach uses a single decoder that takes the concatenated tokens from all modalities as input.
@@
In the literature, MAEs are typically trained using the L1 or L2 losses, which can negatively affect the reconstruction quality of the masked tokens, resulting, for instance, in blurred or noisy images or sounds. It has been shown that improving the quality of MAE reconstructions can be beneficial in terms of downstream task performance [He et al., 2022]. Several approaches have been proposed in that sense, which include adding a perceptual loss [Dong et al., 2021] or using discrete representations obtained from VQ generative adversarial networks (VQ-GANs) [Esser et al., 2021] or variational autoencoders (VQ-VAEs) [Van Den Oord et al., 2017] to train the MAE [Li et al., 2023, Sadok et al., 2023]. These works only considered an unimodal setting, while the present paper proposes a multimodal MAE for audiovisual speech representation learning. Recently, a model called CAV-MAE [Gong et al., 2022b] was proposed combining MAEs and contrastive learning to learn a representation from audiovisual data. However, different from the proposed VQ-MAE-AV model, CAV-MAE processes raw audiovisual data instead of vector-quantized latent representations, it uses simple mean pooling strategies instead of attention pooling, and it is not applied to SER.
@@ 3 The VQ-MAE-AV model
This section presents the VQ-MAE-AV model. The overall approach is illustrated in Fig. 1 and 2 and it can be summarized as follows:
@@
    After self-supervised learning, the VQ-MAE-AV encoder and attention pooling layers are fine-tuned for supervised audiovisual SER (see Section 3.7).
@@ 3.1 Vector quantized variational autoencoder
The proposed multimodal self-supervised approach uses the discrete latent representation of two pre-trained and frozen VQ-VAEs [Van Den Oord et al., 2017]. Specifically, as illustrated in Fig. 1, we use the VQ-VAE-audio and VQ-VAE-visual encoders to obtain compressed and quantized representations of the input speech power spectrogram ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}\in\mathbb{R}^{T_{a}\times D}$ where $T_{a}$ and $D$ correspond to the time and frequency dimensions, and of the input image sequence ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}\in\mathbb{R}^{T_{v}\times H\times W\times C}$ where $T_{v}$, $H$, $W$ and $C$ correspond to the time, height, width, and channel dimensions. The audio and visual quantized representations are denoted by ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$, respectively. Each entry of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ corresponds to the index of a vector in the VQ-VAE codebooks. Notably, ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ retains the time-frequency structure of the original spectrogram, while ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ retains the spatio-temporal structure of the original sequence images. This is because the VQ-VAE-audio and VQ-VAE-visual models are designed to be fully convolutional on the frequency and spatial axes, respectively, and they process the frames within a sequence independently. Therefore, compression occurs along the frequency axis ($D^{\prime}\ll D$) for ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and along the x and y-axes of the image ($H^{\prime}\ll H$, $W^{\prime}\ll W$) for ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$. As shown in Fig. 2 and discussed in the following subsections, the proposed MAE-based self-supervised learning approach operates on these discrete and compressed representations before audiovisual speech reconstruction using the VQ-VAE decoders.
The training procedure of the VQ-VAEs follows the original approach presented in Van Den Oord et al. [2017]. In particular, the VQ-VAE loss functions involve a reconstruction term between the original and reconstructed data, which corresponds to the mean squared error for the visual modality and to the Itakura-Saito divergence for the audio modality [Févotte et al., 2009]. More details are provided in Section 4.1.2.
@@ 3.2 Discrete audio and visual tokens
As shown in Fig. 1, the audio and visual quantized representations ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{T_{v}\times H^{\prime}\times W^{\prime}}$ from the output of the VQ-VAE encoders are divided into non-overlapping patches to build discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot t_{a})\times(n_{d}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot t_{v})\times(n_{h}\cdot h)\times(n_{w}\cdot w)}$, where $T_{a}=n_{t_{a}}\cdot t_{a}$, $D^{\prime}=n_{d}\cdot d$, $T_{v}=n_{t_{v}}\cdot t_{v}$, $H^{\prime}=n_{h}\cdot h$, and $W^{\prime}=n_{w}\cdot w$. These representations are reshaped to ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{Z}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{Z}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$, which are seen as sequences of $n_{t_{a}}\cdot n_{d}$ and $n_{t_{v}}\cdot n_{h}\cdot n_{w}$ tokens of dimension $t_{a}\cdot d$ and $t_{v}\cdot h\cdot w$, respectively.
The use of discrete audio and visual tokens in VQ-MAE-AV has several motivations. Firstly, by dividing the audiovisual data into spatio-spectro-temporal patches, the method could learn to relate audio tokens to visual tokens. For example, the audio tokens are expected to correlate strongly with the visual tokens corresponding to the mouth area [Arnela et al., 2016, Sadok et al., 2024]. The proposed method can potentially learn a representation that captures shared and distinctive information between the two modalities by exploiting their complementarity. Additionally, the use of discrete tokens can reduce the computational cost of the method as it involves working with a reduced representation of the data, which allows us to increase the number of tokens (i.e., manipulate longer audiovisual speech sequences) without exploding in the number of trainable parameters compared to the multimodal MAE in the literature [Bachmann et al., 2022].

### v3

@@ Abstract
An important challenge in emotion recognition is to develop methods that can leverage unlabeled training data. In this paper, we propose the VQ-MAE-AV model, a self-supervised multimodal model that leverages masked autoencoders to learn representations of audiovisual speech without labels. The model includes vector quantized variational autoencoders that compress raw audio and visual speech data into discrete tokens. The audiovisual speech tokens are used to train a multimodal masked autoencoder that consists of an encoder-decoder architecture with attention mechanisms. The model is designed to extract both local (i.e., at the frame level) and global (i.e., at the sequence level) representations of audiovisual speech. During self-supervised pre-training, the VQ-MAE-AV model is trained on a large-scale unlabeled dataset of audiovisual speech, for the task of reconstructing randomly masked audiovisual speech tokens and with a contrastive learning strategy. During this pre-training, the encoder learns to extract a representation of audiovisual speech that can be subsequently leveraged for emotion recognition. During the supervised fine-tuning stage, a small classification model is trained on top of the VQ-MAE-AV encoder for an emotion recognition task. The proposed approach achieves state-of-the-art emotion recognition results across several datasets in both controlled and in-the-wild conditions.
Keywords: Self-supervised learning; Masked autoencoder; Audiovisual speech representation learning; Emotion recognition.
@@
Self-supervised learning , masked autoencoder , audiovisual speech representation learning , emotion recognition.
journal: Computer Vision and Image Understanding
@@ 1 Introduction
Emotions are primarily communicated through nonverbal cues, such as tone, voice, facial expressions, and body language, rather than the words we use in oral communication (Mehrabian, 2017). With the increasing prevalence of audiovisual interfaces, there is a growing demand for systems capable of accurately recognizing emotions from audiovisual speech signals. Indeed, combining audio and visual modalities has proven to be highly effective in various tasks, in particular speech emotion recognition (SER) (El Ayadi et al., 2011; Gao and Grauman, 2019; Zhao et al., 2019; Sadok et al., 2024; Ramachandram and Taylor, 2017; Tsai et al., 2019; Schoneveld et al., 2021).
Supervised learning using labeled datasets is the dominant machine learning paradigm in emotion recognition. However, obtaining datasets with emotion labels is often resource-intensive and impractical to scale. In fact, many emotion recognition datasets rely on actors simulating emotions, which demands significant time and effort during data collection. Another strategy consists of manually collecting and annotating in-the-wild data. However, emotion labeling is an ambiguous task, and annotators may not reach a consensus (Busso et al., 2008).
Therefore, one important challenge is to develop methods that can leverage unlabeled training data. Self-supervised learning (SSL) has recently emerged as a promising technique to address this challenge (Wang et al., 2021; Dib et al., 2023; Chen and Rudnicky, 2023; Liu et al., 2022; Zhang et al., 2022). SSL models are pre-trained on a large-scale unlabeled dataset to solve a pretext task, and they are subsequently fine-tuned on a small amount of labeled data to solve a given task (Gong et al., 2022a; Pepino et al., 2021; Jegorova et al., 2023).
In this paper, we present the first multimodal SSL approach based on masked autoencoders (He et al., 2022) for emotion recognition. We propose the VQ-MAE-AV model, a vector quantized (VQ) masked autoencoder (MAE) designed for audiovisual (AV) speech representation learning and applied to emotion recognition. The proposed approach first involves quantizing the audio and visual speech modalities to create discrete tokens, using vector quantized variational autoencoders (VQ-VAEs). Then, we pre-train the VQ-MAE-AV model in a self-supervised manner to reconstruct audiovisual speech tokens from partially visible inputs, and with an additional contrastive learning strategy. The VQ-MAE-AV model is based on an encoder-decoder architecture with attention mechanisms to fuse the modalities. Its self-supervised pre-training leverages 1 000 hours of audiovisual speech without needing emotion labels, allowing the VQ-MAE-AV model to learn an internal representation of audiovisual speech. Finally, this learned representation is used as input for an emotion recognition model trained on small-scale audiovisual speech datasets labeled with emotion categories. The experimental results demonstrate that the proposed VQ-MAE-AV model achieves superior performance compared to state-of-the-art methods across several datasets in both controlled and in-the-wild conditions. Additionally, extensive ablation experiments are presented to investigate the impact of different model designs.
The paper is organized as follows. In Section 2 we present the related work. The proposed VQ-MAE-AV model is introduced in Section 3. Experiments are presented in Section 4 and we conclude in Section 5. The code and qualitative results are available at https://samsad35.github.io/VQ-MAE-AV.
@@
Fig. 1: Discrete audio and visual tokens creation: (i) fully-convolutional VQ-VAEs are trained independently on the audio and visual modalities (see Section 3.1); (ii) discrete audio and visual tokens are built from the quantized representations provided by the frozen VQ-VAE encoders (see Section 3.2).
@@ 2 Related work
SSL models can be broadly classified into two main categories: discriminative and generative approaches (Zhang et al., 2022). Discriminative SSL focuses on creating pairs or groups of data samples and formulating loss functions that allow the model to differentiate or group these samples, which can later benefit downstream tasks (Chen et al., 2020b, a). For instance, pretext tasks can consist of solving jigsaw puzzles (Noroozi and Favaro, 2016) or predicting image rotations (Gidaris et al., 2018). Contrastive learning has emerged as the predominant paradigm in discriminative SSL (Alayrac et al., 2020; Akbari et al., 2021). In contrast, generative SSL involves generating or reconstructing segments of unlabeled and potentially corrupted data using an encoder-decoder model (He et al., 2022; Bao et al., 2021; Xie et al., 2022). The latent representation produced by the encoder can then be leveraged for downstream tasks.
@@
The present paper focuses on the MAE, an SSL generative model that employs an asymmetric encoder-decoder architecture with input masking (He et al., 2022). The MAE approach is inspired by the concept of masked language modeling (Devlin et al., 2019) and has been successfully applied to image modeling thanks to the development of the Vision Transformer (ViT) (Dosovitskiy et al., 2020). The MAE has recently been extended to audio using a 2D time-frequency representation (Gong et al., 2022a; Baade et al., 2022; Huang et al., 2022). In the MAE paradigm, the input is divided into non-overlapping patches, each represented by a token embedding. Some tokens are masked, typically 75% for image/audio modeling and 15% for text modeling, and only the visible tokens are fed to the encoder. The encoder outputs are used to reconstruct the masked tokens through a lightweight decoder. To successfully reconstruct the masked tokens, the encoder must capture a semantic representation, which can then be effectively transferred to downstream tasks (He et al., 2022). It was recently shown that combining the task of reconstructing masked tokens with contrastive learning can improve the representation learned by an MAE (Huang et al., 2023; Gong et al., 2022a).
Extensions of the MAE for sequential and multimodal representation learning
A recent extension of the MAE was presented in (Feichtenhofer et al., 2022; Tong et al., 2022) for modeling image sequences, called Video-MAE. This method employs the same architecture as the vanilla MAE (He et al., 2022) but incorporates a masking process from video ViT (ViViT) (Arnab et al., 2021). Since videos often contain redundant information, particularly in scenes with no motion, the authors proposed a cubic masking approach along the temporal dimension, combined with a high masking ratio of 90%. Other works have extended the MAE to handle multimodal data (Bachmann et al., 2022; Geng et al., 2022; Gong et al., 2022b). MultiMAE (Bachmann et al., 2022) encodes a small random subset of visible tokens from multiple modalities (RGB, depth, and semantic images) and trains the model to reconstruct the missing tokens. M3AE (Geng et al., 2022) is a unified MAE architecture for two input modalities (image and text). The main difference between M3AE and MultiMAE lies in the architecture of the decoder. The MulitMAE approach introduces individual decoders for each modality, enhancing their fusion by incorporating a cross-attention layer at the beginning of each decoder. On the other hand, the M3AE approach uses a single decoder that takes the concatenated tokens from all modalities as input.
@@
In the literature, MAEs are typically trained using the L1 or L2 losses, which can negatively affect the reconstruction quality of the masked tokens, resulting, for instance, in blurred or noisy images or sounds. Studies have demonstrated that improving the quality of MAE reconstructions can be beneficial in terms of downstream task performance (He et al., 2022). Several approaches have been proposed for that purpose, for instance adding a perceptual loss (Dong et al., 2021) or using discrete representations obtained from VQ generative adversarial networks (VQ-GANs) (Esser et al., 2021) or variational autoencoders (VQ-VAEs) (Van Den Oord et al., 2017) to train the MAE (Li et al., 2023; Sadok et al., 2023). While these works only considered a unimodal setting, the present paper proposes a multimodal MAE for audiovisual speech representation learning. Recently, a model called CAV-MAE (Gong et al., 2022b) was proposed combining MAEs and contrastive learning to learn a representation from audiovisual data. However, unlike the proposed VQ-MAE-AV model, CAV-MAE processes raw audiovisual data instead of vector-quantized latent representations, employs simple mean pooling strategies rather than attention pooling, and has not been applied to SER.
Fig. 2: VQ-MAE-AV model structure. See the first paragraph of Section 3 for a complete description of the pipeline.
@@ 3 The VQ-MAE-AV model
This section presents the VQ-MAE-AV model, which is illustrated in Fig. 1 and 2 and can be summarized as follows:
@@
    After self-supervised learning, a small classification model is trained and the VQ-MAE-AV encoder is fine-tuned for supervised audiovisual SER (see Section 3.7).
@@ 3.1 Vector quantized variational autoencoder
The proposed multimodal self-supervised approach uses the discrete latent representation of two pre-trained and frozen VQ-VAEs (Van Den Oord et al., 2017). Specifically, as illustrated in Fig. 1, we use the VQ-VAE-A (A for audio) and VQ-VAE-V (V for visual) encoders to obtain compressed and quantized representations of the input speech power spectrogram ${\color[rgb]{1,0,0}\mathbf{x}^{(a)}}\in\mathbb{R}^{T_{a}\times D}$ and of the input image sequence ${\color[rgb]{0,0,1}\mathbf{x}^{(v)}}\in\mathbb{R}^{T_{v}\times H\times W\times C}$, where $T_{a}$ and $D$ correspond to the time and frequency dimensions of the audio modality, and $T_{v}$, $H$, $W$ and $C$ correspond to the time, height, width, and channel dimensions of the audio modality. The audio and visual quantized representations are denoted by ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{T_{v}\times H^{\prime}\times W^{\prime}}$, respectively. Each entry of ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ corresponds to the index of a vector in the VQ-VAE codebooks. Notably, ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ retains the time-frequency structure of the original spectrogram, while ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$ retains the spatio-temporal structure of the original sequence images. This is because the VQ-VAE-A and VQ-VAE-V models are designed to be fully convolutional on the frequency and spatial axes, respectively, and they process the frames within a sequence independently. Therefore, compression occurs along the frequency axis ($D^{\prime}\ll D$) for ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}$ and along the x and y-axes of the image ($H^{\prime}\ll H$, $W^{\prime}\ll W$) for ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}$. As shown in Fig. 2 and discussed in the following subsections, the proposed MAE-based self-supervised learning approach operates on these discrete and compressed representations before audiovisual speech reconstruction using the VQ-VAE decoders.
The training procedure of the VQ-VAEs follows the original approach presented in Van Den Oord et al. (2017). In particular, the VQ-VAE loss functions involve a reconstruction term between the original and reconstructed data, which corresponds to the mean squared error for the visual modality and to the Itakura-Saito divergence for the audio modality (Févotte et al., 2009). More details are provided in Section 4.1.2.
@@ 3.2 Discrete audio and visual tokens
As shown in Fig. 1, the audio and visual quantized representations ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{T_{a}\times D^{\prime}}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{T_{v}\times H^{\prime}\times W^{\prime}}$ from the output of the VQ-VAE encoders are divided into non-overlapping patches to build discrete tokens ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot t_{a})\times(n_{d}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{(n_{t_{v}}\cdot t_{v})\times(n_{h}\cdot h)\times(n_{w}\cdot w)}$, where $T_{a}=n_{t_{a}}\cdot t_{a}$, $D^{\prime}=n_{d}\cdot d$, $T_{v}=n_{t_{v}}\cdot t_{v}$, $H^{\prime}=n_{h}\cdot h$, and $W^{\prime}=n_{w}\cdot w$. These representations are reshaped to ${\color[rgb]{1,0,0}\mathbf{x}_{q}^{(a)}}\in\mathbb{N}^{(n_{t_{a}}\cdot n_{d})\times(t_{a}\cdot d)}$ and ${\color[rgb]{0,0,1}\mathbf{x}_{q}^{(v)}}\in\mathbb{N}^{(n_{t_{v}}\cdot n_{h}\cdot n_{w})\times(t_{v}\cdot h\cdot w)}$, which are seen as sequences of $n_{t_{a}}\cdot n_{d}$ and $n_{t_{v}}\cdot n_{h}\cdot n_{w}$ tokens of dimension $t_{a}\cdot d$ and $t_{v}\cdot h\cdot w$, respectively.
The use of discrete audio and visual tokens in VQ-MAE-AV has several motivations. Firstly, by dividing the audiovisual data into spatio-spectro-temporal patches, the method could learn to relate audio tokens to visual tokens. For example, the audio tokens are expected to correlate strongly with the visual tokens corresponding to the mouth area (Arnela et al., 2016; Sadok et al., 2024). The proposed method can potentially learn a representation that captures shared and distinctive information between the two modalities by exploiting their complementarity. Additionally, the use of discrete tokens can reduce the computational cost of the method as it involves working with a reduced representation of the data, which allows us to increase the number of tokens (i.e., manipulate longer audiovisual speech sequences) without exploding in the number of trainable parameters compared to the multimodal MAE in the literature (Bachmann et al., 2022).